UNRAVELING THE COMPLEX NEAR-UV SPECTRUM OF Si_{2}C

\underline{N. J. REILLY} and M. C. McCARTHY, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; D. L. KOKKIN, IRAP; Université de Toulouse, UPS; CNRS; 9 Av. colonel Roche, BP 44346, F-31028 Toulouse cedex 4, France; R. C. FORTENBERRY, and T. D. CRAWFORD, Department of Chemistry, Virginia Tech, Blacksburg, VA 24061.

Though it has yet to be detected in space, there is good reason to believe that the disilicon carbide molecule, Si_{2}C, is an abundant constituent of circumstellar shells such as IRC+10216, in which the isovalent ionic ring SiC_{2} is highly conspicuous. Si_{2}C exhibits a surprisingly complicated R2C2PI spectrum, which EOM-CCSD calculations indicate arises from a bent-to-linear transition involving several interacting electronic states in the region of vertical excitation. Some of the observed spectral features show well-resolved emission spectra when probed by LIF/DF, yielding vibrational frequencies in excellent agreement with ab initio and matrix IR studies of the ground electronic state. Partially resolved rotational structure for these bands, which appear to be dominated by a vibronically allowed transition to a dark A_{2} state, suggests a ground state bond angle significantly larger than predicted by theory. Owing to the presence of overlapping, unresolved bands, in addition to spectral contamination from Si_{2} and SiC, further experimental characterization would appear ideally suited to a two-dimensional fluorescence approach.